A mortgage broker is offering a 25-year $270,000 mortgage with a teaser rate. In the first two years of the mortgage, the borrower makes monthly payments on only a 3.6 percent APR interest rate. After the second year, the mortgage interest rate charged increases to 6.6 percent APR.

What are the monthly payments in the first two years?

What are the monthly payments after the second year?
Answer #1

Using Microsoft Excel
1
=PMT(3.6%/12,12*25,-270000)=1366.20732358758

2
=PMT(6.6%/12,12*24,FV(3.6%/12,12,-PMT(3.6%/12,12*25,-270000),270000))=1823.37639050531